OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of 2 and a y-intercept of -3.

OpenStudy (♪chibiterasu):

All we have to do is replace the variables in y = mx + b with the corresponding number. y is y-intercept and m is the slope. We have a y-intercept of -3 so we can plug it into y, and our slope is 2 so we can plug it into m. y = mx + b -3 = 2x + b

OpenStudy (♪chibiterasu):

Pardon the y-intercept is b

OpenStudy (♪chibiterasu):

y = 2x -3

OpenStudy (♪chibiterasu):

Your welcome >.>

OpenStudy (anonymous):

thank you :D

OpenStudy (♪chibiterasu):

No problem :3